OpenStudy (anonymous):

In 2010 a company installed a new machine in one of its factories at a cost of $680,000. The machine is linearly depreciated over 30 years with a scrap value of $80,000. 4. Find an expression for the machine value in the t th year of use (0 < t < 30).

OpenStudy (danjs):

sorry, it is scrap value...

OpenStudy (danjs):

linearly depreciated means it is a line with constant slope, y = m*x+b

OpenStudy (anonymous):

I got -20000 so m ans is v(t) = -20,000t + 680000?

OpenStudy (danjs):

the slope will be how much it has gone down in over that 30 years time... m = (680,000 - 80,000) / 30

OpenStudy (danjs):

looks good, the y intercept is the starting value when time t=0

OpenStudy (anonymous):

my answer is right? the second part is: Find the machine’s value in the year 2024. how do I calculate that?

OpenStudy (danjs):

yes correct, they installed the achine in 2010, so 2024-2010 = 14 years of depreciation, let t=14 years